# Mosaic3D: Foundation Dataset and Model for Open-Vocabulary 3D Segmentation

### v1

@@ Abstract
[…]
Figure 1: Mosaic3D-5.6M. Mosaic3D-5.6M is a large-scale dataset generated from ScanNet [22], ARKitScenes [5], ScanNet++ [93], Matterport3D [12], and Structured3D [101], consisting of 5.6M mask-text pairs, providing fine-grained masks (black outline in the figure) and detailed captions (text with matching color) pairs. Using this large-scale dataset, we propose Mosaic3D, a foundation model for open-vocabulary 3D segmentation.
@@ 1 Introduction
3D scene understanding is a fundamental problem in computer vision that involves detecting and localizing objects while comprehending the complex spatial relationships in 3D environments. This capability is essential for diverse applications, including robotics, AR/VR, human-computer interactions, and autonomous vehicles. While traditional approaches rely on predefined object categories, the field is evolving toward open-vocabulary 3D scene understanding—where systems can recognize arbitrary concepts without being constrained to the predefined label sets. Despite humans’ innate ability to perform such tasks, developing comparable machine capabilities remains an open problem.
@@
The key bottleneck in advancing open-vocabulary 3D scene understanding is the scarcity of large-scale, high-quality training data. This limitation is particularly striking when compared to 2D vision-language models [66, 36, 48, 29, 98, 52, 50, 51, 32, 53, 68, 77, 89], which have achieved remarkable open-vocabulary capabilities through training on web-scale image-text pair datasets [66, 75, 9, 14, 28]. Unfortunately, creating datasets of comparable scale for 3D scenes remains prohibitively expensive and time-consuming.
@@
Training effective open-vocabulary 3D scene understanding models requires datasets that satisfy three critical requirements: (1) precise segmentation of 3D regions that accurately delineate object boundaries, (2) comprehensive textual descriptions that capture the detailed characteristics of each region, and (3) sufficient dataset size with many annotated scenes. Creating such data manually, however, becomes increasingly intractable as datasets grow.
@@
To address these challenges, recent works [25, 92, 45] leverage 2D Visual Foundation Models (VFMs) [55, 1, 64, 104] to automate data annotation. They generate 3D mask-text pairs on multi-view RGB-D frames using 2D VFMs and later aggregate the generated captions in 3D space. However, they fall short in meeting aforementioned requirements: they either use coarse bounding box detectors [25, 92] that cannot capture precise object boundaries or rely on simple attribute labels [45] instead of detailed descriptions. Additionally, existing datasets are limited in scale, containing only a few thousand scenes, as shown in Fig 4.
@@
In this paper, we address these limitations by introducing an improved data generation pipeline for creating large-scale, high-quality 3D mask-text pairs. Our pipeline satisfies all three criteria by leveraging state-of-the-art open-vocabulary image segmentation models [56, 47, 70, 104] for precise region segmentation and advanced region-aware Vision-Language Models (VLMs) [97, 95] for generating comprehensive textual descriptions at scale. By applying this pipeline to a diverse collection of 3D scene datasets [22, 93, 5, 12, 101], we create Mosaic3D-5.6M, a large dataset containing over 30K scenes with 5.6M region captions—significantly surpassing existing datasets [92, 45] in scale—while maintaining high-quality region masks and detailed textual descriptions. With this expanded training dataset, our method strengthens 3D-language alignment and advances the state-of-the-art in open-vocabulary 3D scene understanding.
@@
Building upon this new dataset, we analyze how scaling up the data and enhancing annotation quality impacts open-vocabulary 3D scene understanding. To enable this analysis, we develop a general framework for open-vocabulary 3D semantic and instance segmentation. We train our foundational 3D encoder, Mosaic3D, based on Sparse ConvNets [31, 19] that aligns per-point features with text embeddings through contrastive learning. Then, we train a lightweight mask decoder to predict object instances directly from language-aligned features, enabling the first single-stage open-vocabulary 3D instance segmentation without ground truth labels. Our approach achieves state-of-the-art results on multiple semantic and instance segmentation benchmarks. Extensive ablation studies reveal that both the scale and the quality of our dataset are essential components, as our dataset yields significantly better performance compared to smaller or lower quality alternatives.
Figure 2: Dataset comparison. Left: Original point cloud. (a) Our Mosaic3D-5.6M leverages precise segmentation masks and region-aware VLMs to generate detailed descriptions. (b) OV3D [45] only generates simple object attribute labels rather than comprehensive captions that reflect visual and geometric observations. (c) RegionPLC [92] relies on coarse object bounding boxes, resulting in imprecise masks.
@@ 2 Related Work
2D Visual Foundation Models. In recent years, we have witnessed the emergence of large pretrained models—so-called Foundation Models (FMs)—that are trained on broad sets of data and serve as the foundation for many downstream deep learning tasks. These models demonstrate remarkable versatility across multiple modalities , including language [82, 84, 85, 26, 17, 65, 8, 20, 2, 3, 91, 43, 44] , vision [47, 70, 11, 62, 104, 72, 35, 61, 79, 80] , audio [24, 99, 74, 6]. Furthermore, they enable multimodal reasoning capabilities that bridge across different modalities [30, 4, 55, 66, 42, 83]. Among these models, those that operate on visual modalities are known as Visual Foundation Models (VFMs).
Visual Foundation Models excel at various computer vision tasks such as image segmentation [47, 70, 104, 103, 15, 16, 40, 49], object detection [56, 10], representation learning [11, 62], and open-vocabulary understanding [66, 48, 29, 100, 38, 96, 46, 58, 18]. When integrated with Large Language Models, they enable sophisticated visual reasoning and natural language interactions [55, 4, 30, 83, 33, 97, 95]. Rapid advancement of these models has been driven by large-scale internet-scraped datasets that span images, videos, and multimodal content [75, 9, 14, 28].
Open-vocabulary 3D semantic segmentation. Building on the success of 2D VFMs, recent work have explored extending open-vocabulary capabilities to 3D scene understanding. OpenScene [63] pioneered this direction by distilling knowledge from language-aligned image encoders [48, 29] into 3D point cloud encoders, enabling zero-shot semantic segmentation of 3D scenes. To address inconsistency from multiview image distillation, subsequent approaches such as PLA [25], RegionPLC [92] and OV3D [45] leverage multiview images to generate textual descriptions, which then serve as training supervision. However, these methods face challenges in generating accurate and descriptive mask-text pairs on a scale.
Open-vocabulary 3D instance segmentation. In parallel with progress on semantic segmentation, recent work has also explored open-vocabulary instance segmentation in 3D scenes. Current methods can be grouped into two categories based on their region proposal approach and inference requirements. The first category [81, 60, 39] depends on closed-vocabulary 3D region proposal networks such as Mask3D [76], which inherently constrains their ability to detect novel object categories. Although SAI3D [94] attempted to address this using training-free superpoints [27], it still faces the second limitation. The second group [81, 60, 39] borrows the powerful open-vocabulary capabilities of 2D VFMs, such as CLIP [66], to classify the detected regions by projecting them onto 2D views. However, they require both 2D images and 3D point clouds during inference, performing multiple inferences of large 2D models on projected masks, making them computationally expensive. Even OpenIns3D [39], which processes 3D point clouds directly, still requires projecting regions to 2D for classification [90]. More recently, Segment3D [37] made progress on the first limitation by achieving class-agnostic 3D region proposals using SAM [47] supervision, but it still requires 2D CLIP inference as its mask embeddings are not text-aligned.
We address these limitations by developing the first single-stage open-vocabulary 3D instance segmentation model that operates directly in 3D without ground truth labels, using our Mosaic3D-5.6M dataset and Segment3D [37] proposals.
@@
As 3D mask-text pair datasets are scarce and expensive to create manually, recent work [25, 92, 45] have leveraged 2D Visual Foundation Models (VFMs) to automate data annotation. These methods utilize multi-view images as a bridge to transfer 2D VFM’s open-vocabulary understanding capabilities into 3D space by a combination of 2D region proposals (e.g. object detection, segmentation, or sliding window) and image captioning. However, existing approaches either suffer from imprecise boundary delineation due to their reliance on coarse object detection models [25, 92], or provide only simple attribute labels instead of comprehensive descriptions [45]. To overcome these limitations, we propose a data generation pipeline that combines recent advances in open-vocabulary segmentation and robust region-aware Vision-Language Models (VLMs), enabling both precise region boundaries and rich descriptions that capture object attributes, spatial relationships, and scene context.
@@ 3.1 Proposed Pipeline
Our pipeline overcomes limitations through two key improvements. Fig. 3 illustrates the overview of our pipeline.
@@
Enhanced segmentation. We employ Grounded-SAM [71] and SEEM [104] for more precise open-vocabulary image segmentation. Unlike previous methods [92, 25] that rely on coarse object detectors, these models enable more accurate boundary delineation and can identify arbitrary objects within the scene. The reason we incorporate both models is that Grounded-SAM excels at segmenting foreground objects with precise boundaries, while SEEM complements this by performing open-vocabulary panoptic segmentation that better handles background stuff like wall and floor.
@@
Given an RGB image $\mathbf{I}$, Grounded-SAM first predicts open-set bounding boxes using Grounding-DINO [56], then uses these boxes as input prompts for SAM [47, 70] to generate segmentation masks. To fully automate the segmentation process, we employ RAM++ [38] to detect object categories within the input image. These detected categories serve as a text prompt for Grounding-DINO. Through this, Grounded-SAM generates a set of segmentation masks $\{\mathbf{M}_{k}\}_{k=1}^{K}$ as:
$\{\mathbf{M}_{k}\}_{k=1}^{K}=\text{Grounded-SAM}(\mathbf{I}),$ (1)
where $K$ is the number of detected objects in the image. Each mask $\mathbf{M}_{k}\in\{0,1\}^{H\times W}$ represents a binary segmentation of an object. SEEM operates in a similar way except that it directly performs panoptic segmentation without being prepended with RAM++ and Grounding-DINO for object detection. For simplicity, we denote the combined set of masks from both Grounded-SAM and SEEM as $\{\mathbf{M}_{k}\}_{k=1}^{K}$, where $K$ is the total number of masks from both models.
@@
Enhanced region captioning. For each segmentation mask, we generate a detailed caption that describes the object’s visual characteristics and spatial context. Unlike previous methods that process holistic images using generic image captioning models [64, 55, 1, 87] , we leverage region-aware Vision-Language Models (VLMs) [95, 97, 69] that are specifically designed to understand and describe localized regions in images. These models can generate detailed descriptions by interpreting various visual prompts like points, boxes, masks, and scribbles, enabling more focused and contextual captions for each segmented region.
@@
Given an image $\mathbf{I}$, a segmentation mask $\mathbf{M}$, and a user prompt $\pi$ that asks for a detailed description of the masked region, the region-aware VLM $\mathcal{R}(\cdot)$ generates a textual response $c$ for each mask by $c=\mathcal{R}(\mathbf{I};\mathbf{M},\pi)$, where $c$ is a natural language description that captures the visual attributes and spatial context of the masked region. After evaluating several available region-aware VLMs, we chose Osprey [97] for our implementation.
@@
2D pixel-3D point association. After obtaining segmentation masks and captions from multiple views, we associate them with 3D points to create mask-text pairs in 3D space. For each 3D point $\mathbf{p}$ in the point cloud $\mathbf{P}\in\mathbb{R}^{N\times 3}$, we project it onto each view using the camera parameters to obtain its 2D pixel coordinates $(u,v)$ and depth value $d$. We then perform an inclusion test by checking if the projected pixel falls within any segmentation mask $\mathbf{M}_{k}$ and if its depth matches the image depth at that location within a small threshold $\epsilon$. Specifically, for point cloud $\mathbf{P}$, we compute 3D binary region masks $\mathbf{s}_{k}\in\{0,1\}^{N}$ as:
@@
$\!\!\mathbf{s}_{k}=\left\{\begin{array}[]{cl}\!1&\!\!\!\textrm{if}~{}~{}\mathbf{M}_{k}(u,v)=1\textrm{ and }|d-\mathbf{D}(u,v)|_{1}<\epsilon\\ \!0&\!\!\!\textrm{otherwise}\end{array}\right.\!\!\!\!,$ (2)
@@
where $\mathbf{D}$ is the depth map corresponding to image $\mathbf{I}$. Finally, we obtain 3D mask-text pairs $\{(\mathbf{s}_{k},c_{k})\}_{k=1}^{K}$ that associate each segmented region with its corresponding caption.
@@
Our pipeline produces high-quality 3D mask-text pairs that combine precise object boundaries with rich semantic descriptions. As demonstrated in Fig. 2, compared to previous methods, our approach achieves both more accurate segmentation boundaries and more detailed, contextual descriptions that capture object attributes and spatial relationships.

### v2

@@ Abstract
We tackle open-vocabulary 3D scene segmentation tasks by introducing a novel data generation pipeline and training framework. Our work targets three essential aspects required for an effective dataset: precise 3D region segmentation, comprehensive textual descriptions, and sufficient dataset scale. By leveraging state-of-the-art open-vocabulary image segmentation models and region-aware vision-language models (VLM), we develop an automatic pipeline capable of producing high-quality 3D mask-text pairs. Applying this pipeline to multiple 3D scene datasets, we create Mosaic3D-5.6M, a dataset of more than 30K annotated scenes with 5.6M mask-text pairs - significantly larger than existing datasets. Building on these data, we propose Mosaic3D, a 3D visiual foundation model (3D-VFM) combining a 3D encoder trained with contrastive learning and a lightweight mask decoder for open-vocabulary 3D semantic and instance segmentation. Our approach achieves state-of-the-art results on open-vocabulary 3D semantic and instance segmentation benchmarks including ScanNet200, Matterport3D, and ScanNet++, with ablation studies validating the effectiveness of our large-scale training data. https://nvlabs.github.io/Mosaic3D/ Authors equally contributed to this work during internship at NVIDIA.
Figure 1: Mosaic3D-5.6M. Mosaic3D-5.6M is a large-scale dataset generated from a collection of existing datasets [24, 7, 99, 14, 107], consisting of 5.6M mask-text pairs, providing fine-grained masks (black outline in the figure) and detailed captions (text with matching color) pairs. Using this large-scale dataset, we propose Mosaic3D, a foundation model for open-vocabulary 3D segmentation.
@@ 1 Introduction
3D scene understanding is a fundamental problem in computer vision that involves detecting and localizing objects while comprehending complex spatial relationships in 3D environments. This capability is essential for various applications, including robotics, AR/VR, human-computer interactions, and autonomous vehicles. While traditional approaches rely on predefined object categories, the field is evolving toward open-vocabulary 3D scene understanding, where systems can recognize arbitrary concepts without being constrained to the predefined label sets. Despite humans’ innate ability to perform such tasks effortlessly, developing comparable machine capabilities remains an open problem.
@@
The key bottleneck in advancing open-vocabulary 3D scene understanding is the scarcity of large-scale, high-quality training data. This limitation is particularly striking compared to 2D vision-language models [71, 38, 51, 31, 104, 55, 53, 54, 34, 56, 73, 83, 96], which have achieved remarkable open-vocabulary capabilities through training on web-scale image-text pair datasets [71, 81, 11, 16, 30]. Unfortunately, creating datasets of comparable scale for 3D scenes remains prohibitively expensive and time-consuming.
@@
Training effective open-vocabulary 3D scene understanding models requires datasets that satisfy three critical requirements: (1) precise 3D region annotations that delineate object boundaries, (2) rich textual captions that characterize the visual and semantic attributes of each region, and (3) substantial scale to encompass diverse domains and offer rich visual-semantic variations. Creating such data manually, however, becomes increasingly intractable as datasets grow.
@@
To address these challenges, recent works [27, 98, 48] leverage 2D visual foundation models (VFM) [58, 1, 69, 110] to automate data annotation. They generate 3D mask-text pairs on multi-view RGB-D frames using 2D VFMs and aggregate the generated captions in 3D space. However, they fall short in meeting aforementioned requirements: they use coarse bounding box detectors [27, 98]; only have simple attribute labels [48]; and are limited in scale, containing only a few thousand scenes, as shown in Fig. 4. Existing 3D-text pair datasets for 3D vision-language models [44, 94, 60] also face limitations in both the richness of textual captions and the precision of 3D masks, primarily because they rely on human-annotated object labels and 3D region annotations.
@@
In this paper, we address these limitations by introducing an improved data generation pipeline to create high-quality, large-scale 3D mask-text pairs. Our pipeline satisfies all three criteria by leveraging state-of-the-art open-vocabulary image segmentation models [59, 50, 75, 110] for precise region segmentation and advanced region-aware vision-language models [103] for generating comprehensive textual captions at scale. By applying this pipeline to a diverse collection of 3D scene datasets [24, 99, 7, 14, 107], we create Mosaic3D-5.6M, a large dataset containing over 30K scenes with 5.6M region captions – significantly exceeding existing datasets [98, 48] in scale – while maintaining high-quality region masks and detailed textual captions.
@@
Building upon this new dataset, we analyze how improving the annotation quality and scaling up the data impact open-vocabulary 3D scene segmentation. To enable this analysis, we develop a general framework for open-vocabulary 3D semantic and instance segmentation. We train our foundational 3D encoder, Mosaic3D, which aligns a per-point feature with a text embedding vector through contrastive learning. Then, we train a lightweight mask decoder to predict object instances directly from language-aligned features, enabling the first single-stage open-vocabulary 3D instance segmentation without ground truth labels. Our approach achieves state-of-the-art results on multiple semantic and instance segmentation benchmarks. Extensive ablation studies show that both the scale and the quality of our dataset are crucial factors that contribute to the superior performance of our approach.
Figure 2: Dataset comparison. We compare datasets using three metrics: # Nouns (the total number of unique normalized nouns in captions; higher is better), Coverage (the percentage of 3D points with associated captions per scene; higher is better), and Entropy (the entropy of GT instance ID distribution within masks; lower means more homogeneity - hense better). (a) Mosaic3D-5.6M uses precise masks (Entropy: 60.7) with region-aware VLMs for detailed descriptions (# Nouns: 29.9K). (b) OV3D [48] produces simple attribute labels (# Nouns: 2.5K) lacking comprehensive visual descriptions. (c) RegionPLC [98] uses coarse bounding boxes, yielding imprecise masks (Entropy: 81.0).
@@ 2 Related Work
2D visual foundation models. In recent years, we have witnessed the emergence of large pretrained models—so-called foundation models that are trained on large-scale datasets and serve as a foundation for many downstream tasks. These models demonstrate remarkable versatility across multiple modalities, including language [88, 90, 91, 28, 19, 70, 10, 22, 2, 5, 97, 46, 47], vision [50, 75, 13, 67, 110, 77, 37, 66, 85, 86], audio [26, 105, 79, 8]. Furthermore, they enable multi-modal reasoning capabilities that bridge across different modalities [32, 6, 58, 71, 45, 89]. Among these models, those that operate on visual modalities are known as visual foundation models (VFM). VFMs excel in various computer vision tasks such as image segmentation [50, 75, 110, 109, 17, 18, 43, 52], object detection [59, 12], representation learning [13, 67], and open-vocabulary understanding [71, 51, 31, 106, 41, 102, 49, 63, 20]. When integrated with large language models, they enable sophisticated visual reasoning and natural language interactions [58, 6, 32, 89, 35, 103, 101]. We use such vision language models to construct open vocabulary segmentation and captions for point clouds based on multiview images.
Open-vocabulary 3D segmentation. Building on the success of 2D VFMs, recent work have extended open-vocabulary capabilities to 3D scene understanding. OpenScene [68] first introduced zero-shot 3D semantic segmentation by distilling knowledge from language-aligned image encoders [51, 31]. Subsequent methods [27, 98, 48] leverage multiview images to generate textual captions, which then serve as training supervision. However, these methods face challenges in generating high-quality 3D mask-text pairs at scale. For open-vocabulary 3D instance segmentation, existing methods [87, 65, 42] typically rely on closed-vocabulary proposal networks such as Mask3D [82], which inherently constrains their ability to detect novel object categories. Moreover, these methods leverage 2D VFMs like CLIP [71] for region classification by projecting 3D regions onto multiple 2D views. This approach requires both 2D images and 3D point clouds during inference. Additionally, it necessitates multiple inferences of large 2D models on projected masks, resulting in high computational costs. We address these limitations by developing the first single-stage open-vocabulary 3D instance segmentation model that operates directly in 3D without ground truth labels, using our Mosaic3D-5.6M dataset and Segment3D [40] proposals.
3D vision-language datasets. Several datasets align 3D scenes with textual annotations to facilitate language-driven 3D understanding. ScanRefer [15], ReferIt3D [3] and EmbodiedScan [94] provide fine-grained object-level localization through detailed referential phrases, while ScanQA [4] targets spatially grounded question-answering. In contrast, SceneVerse [44] and MMScan [60] employ large-language models or vision-language models to partially automate annotation. Despite leveraging advanced models, these datasets depend significantly on costly human annotations derived from closed-vocabulary sources, limiting their support for open-vocabulary and scalability for large-scale 3D segmentation tasks.
@@
Generating 3D mask-text pair datasets can be costly and require meticulous attention. Recent work [27, 98, 48] have leveraged 2D visual foundation models (VFMs) to automate data annotation to an extent – they use multi-view images to generate captions or features on different types of region proposals (e.g. bounding boxes, segmentations, or sliding windows). However, existing approaches suffer from imprecise boundary delineation due to their reliance on coarse object detectors [27, 98], or provide only simple attribute labels [48]. To overcome these limitations, we propose a data generation pipeline that combines recent advances in open-vocabulary segmentation and robust region-aware vision-language models (VLMs), enabling both precise region boundaries and rich descriptions that capture object attributes, spatial relationships, and scene context.
@@ 3.1 Proposed Pipeline
Our pipeline overcomes limitations through two key improvements: accurate segmentation and the region captioning pipeline. Fig. 3 illustrates the overview of our pipeline.
@@
Enhanced segmentation. We employ Grounded-SAM [76] and SEEM [110] for more precise open-vocabulary image segmentation. We incorporate both models because Grounded-SAM excels at segmenting foreground objects with precise boundaries, while SEEM complements this with open-vocabulary panoptic segmentation that better handles background stuff like wall and floor. Thus, our method outperforms previous methods [98, 27] that rely on foreground object detectors.
@@
Given an RGB image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, Grounded-SAM first predicts open-set bounding boxes using Grounding-DINO [59], then uses these boxes as input prompts for SAM [50, 75] to generate segmentation masks. To fully automate the segmentation process, we employ RAM++ [41] to detect object categories within the input image. These detected categories serve as a text prompt for Grounding-DINO. Through this, Grounded-SAM generates a set of segmentation masks $\{\mathbf{M}_{k}\}_{k=1}^{K}$ where $K$ is the number of detected objects in the image. Each mask $\mathbf{M}_{k}\in\{0,1\}^{H\times W}$ represents a binary segmentation of an object. SEEM operates in a similar way, except that it directly performs panoptic segmentation without being required to use RAM++ to generate tags or Grounding-DINO for object detection. For simplicity, we denote the combined set of masks from both Grounded-SAM and SEEM as $\{\mathbf{M}_{k}\}_{k=1}^{K}$, where $K$ is the total number of masks from both models.
@@
Enhanced region captioning. For each segmentation mask, we generate a detailed caption that describes the visual characteristics and spatial context of the object. Unlike previous methods that process an image as a whole using generic image captioning models [69, 58, 1, 93], we leverage region-aware vision-language models (VLMs) [101, 103, 74] that are specifically designed to understand and describe a region or a mask specified by a user as an additional. These models can generate detailed descriptions by interpreting various visual prompts, such as points, boxes, masks, and scribbles, enabling more focused and contextual captions for each segmented region.
@@
Given an image $\mathbf{I}$, a segmentation mask $\mathbf{M}_{k}$, and a user prompt $\pi$ that asks for a detailed description of the masked region, the region-aware VLM $\mathcal{R}(\cdot)$ generates a textual response $c$ for each mask by $c_{k}=\mathcal{R}(\mathbf{I};\mathbf{M}_{k},\pi)$, where $c_{k}$ is a natural language description that captures the visual attributes and spatial context of the k-th masked region. After evaluating several available region-aware VLMs, we chose Osprey [103] for our implementation.
@@
2D pixel-3D point association. After obtaining segmentation masks and captions from multiple views, we associate them with 3D points to create mask-text pairs in 3D space. For each 3D point $\mathbf{p}$ in the point cloud $\mathbf{P}\in\mathbb{R}^{N\times 3}$, we project it onto each view using the camera parameters to obtain its 2D pixel coordinates $(u,v)$ and depth value $d$. We check if the projected pixel falls within any segmentation mask $\mathbf{M}_{k}$ and if its depth matches the ground-truth depth at that location within a small threshold $\epsilon$ (inclusion test). Specifically, for point cloud $\mathbf{P}$, we compute 3D binary region masks $\mathbf{s}_{k}\in\{0,1\}^{N}$ as:
@@
$\!\!\mathbf{s}_{k}=\left\{\begin{array}[]{cl}\!1&\!\!\!\textrm{if}~{}~{}(\mathbf{M}_{k})_{u,v}=1\textrm{ and }|d-\mathbf{D}_{u,v}|_{1}<\epsilon\\ \!0&\!\!\!\textrm{otherwise}\end{array}\right.\!\!\!\!,$ (1)
@@
where $\mathbf{D}$ is the ground-truth depth image. Finally, we obtain 3D mask-text pairs $\{(\mathbf{s}_{k},c_{k})\}_{k=1}^{K}$ that associate each segmented region with its corresponding caption.
@@
Our pipeline produces high-quality 3D mask-text pairs that combine precise object boundaries with rich semantic descriptions. As demonstrated in Fig. 2, compared to previous methods, our approach achieves both more accurate segmentation boundaries and more detailed contextual descriptions that capture object attributes and spatial relationships.